# Singular dynamics for discrete weak K.A.M. solutions of exact twist maps

### v1

@@ 2. Preliminary
(1). To show that $F^{-n}\left(\overline{\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)}\right)\subset\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)$ for all $n\geq 1,$ it suffices to demonstrate
$F^{-1}\left(\overline{\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)}\right)\subset\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u).$
Consider a point $(y,p^{\prime})\in\overline{\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)}$. There exists a sequence $\{(y_{k},p^{\prime}_{k})\}_{k\in\mathbb{Z}_{+}}\subset\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)$ that converges to $(y,p^{\prime})$ as $k\rightarrow+\infty$. If $(y_{k},p^{\prime}_{k})\in\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)$, then $u$ is differentiable at $y_{k}$ and $p^{\prime}_{k}=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y_{k})$. Since $u=T^{-}[u]-\bar{S}$, the function $T^{-}[u]$ is also differentiable at $y_{k}$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{-}[u](y_{k})=p^{\prime}_{k}$. Using item (a) of Lemma 2.6, there exists a unique $x_{k}\in\mathbb{R}^{d}$ such that $T^{-}[u](y_{k})=u(x_{k})+S(x_{k},y_{k})$. By item (a) of Lemma 2.5, we have
$\partial_{2}S(x_{k},y_{k})=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{-}[u](y_{k})=p^{\prime}_{k}.$
Since $S$ is is ferromagnetic, the convergence $(y_{k},p_{k}^{\prime})\rightarrow(y,p^{\prime})$ implies that $\{x_{k}\}_{k\in\mathbb{Z}_{+}}$ converges to some $x\in\mathbb{R}^{d}$ and
$\partial_{2}S(x,y)=p^{\prime}.$
Then, from $T^{-}[u](y_{k})=u(x_{k})+S(x_{k},y_{k})$, we get
$T^{-}[u](y)=u(x)+S(x,y).$
Using $u=T^{-}[u]-\bar{S}$ again, we have
$u(y)+\bar{S}=T^{-}[u](y)=u(x)+S(x,y).$
That is, $(x,y)\in\mathcal{NS}(u)=\mathcal{NS}(T^{-}[u])$. Then it follows from the item (a) of Lemma 2.12 that $u$ is differentiable at $x$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)=-\partial_{1}S(x,y)$. Combine $\partial_{2}S(x,y)=p^{\prime}$ and $-\partial_{1}S(x,y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)$, by Remark 2.3, we have
@@
$F^{-1}(y,p^{\prime})=(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)).$
@@
Therefore, $F^{-1}(y,p^{\prime})\in\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)$.
(2). Let $\widetilde{u}$ be a $C^{1,1}$ sub-action such that $\mathcal{A}=\pi_{1}(\mathcal{NS}(\widetilde{u}))=\pi_{2}(\mathcal{NS}(\widetilde{u}))$ in [Zav10]. Thanks to Remark 2.9, we can define
$\widetilde{u}_{-}:=\lim_{n\to+\infty}T^{-n}[\widetilde{u}]-n\bar{S}.$
Consider a point $(x_{0},p_{0})\in\mathcal{I}^{*}(\widetilde{u}_{-})$ and denote $(x_{k},p_{k}):=F^{k}(x_{0},p_{0}),\ \forall k\in\mathbb{Z}_{+}$. Then we know that $\widetilde{u}_{-}$ is differentiable at $x_{k}$ with differential $p_{k}$ for each $k\geq 1$. Using item (a) of Lemma 2.6, there exists a unique $w_{k}\in\mathbb{R}^{d}$ such that $T^{-}[\widetilde{u}_{-}](x_{k})=\widetilde{u}_{-}(w_{k})+S(w_{k},x_{k})$ for each $k\geq 1$. By item (a) of Lemma 2.5, we have
$\partial_{2}S(w_{k},x_{k})=p_{k},\quad\forall k\geq 1.$
Since $S$ is ferromagnetic and $F(x_{k-1},p_{k-1})=(x_{k},p_{k})$, we have $x_{k-1}=w_{k}$ for each $k\geq 1$. In particular,
$\widetilde{u}_{-}(x_{k})+\bar{S}=T^{-}[\widetilde{u}_{-}](x_{k})=\widetilde{u}_{-}(x_{k-1})+S(x_{k-1},x_{k}),\quad\forall k\geq 1.$
That is, the pair $(x_{k-1},x_{k})$ is $\widetilde{u}_{-}$-calibrated for each $k\geq 1$. Then, by Remark 2.11,
$T^{+}[\widetilde{u}_{-}](x_{k})-T^{+}[\widetilde{u}_{-}](x_{k-1})$
$=$ $\left(\widetilde{u}_{-}(x_{k})-\bar{S}\right)-\left(\widetilde{u}_{-}(x_{k})-S(x_{k-1},x_{k})\right)$
$=$ $S(x_{k-1},x_{k})-\bar{S}.$
That is, the pair $(x_{k-1},x_{k})$ is $T^{+}[\widetilde{u}_{-}]$-calibrated for each $k\geq 1$. We claim, by induction, that for all $k,n\in\mathbb{Z}_{+}$, the pair $(x_{k},x_{k+1})$ is $T^{+n}[\widetilde{u}_{-}]$-calibrated. Specifically, $(x_{0},x_{1})$ is $T^{+n}[\widetilde{u}_{-}]$-calibrated for each $n\in\mathbb{Z}_{+}$. Utilizing Remark 2.11 once again, we obtain
$T^{+n}[\widetilde{u}_{-}](x_{0})+n\bar{S}=\widetilde{u}_{-}(x_{0}),\quad\forall n\in\mathbb{Z}_{+}.$
Let $\varepsilon(n):=\sup\left(\widetilde{u}_{-}-T^{-n}[\widetilde{u}]+n\bar{S}\right)$. Then, based on item (b) of Remark 2.4, for each $n\in\mathbb{Z}_{+}$, we have
$\widetilde{u}_{-}-\widetilde{u}$ $\leq\widetilde{u}_{-}-T^{+n}\circ T^{-n}[\widetilde{u}]$
$\leq\widetilde{u}_{-}-T^{+n}[\widetilde{u}_{-}+n\bar{S}-\varepsilon(n)]$
$=\left(\widetilde{u}_{-}-T^{+n}[\widetilde{u}_{-}]-n\bar{S}\right)+\varepsilon(n).$
Setting $x=x_{0}$ and letting $n\to+\infty$, we obtain $\widetilde{u}_{-}(x_{0})-\widetilde{u}(x_{0})\leq 0$. Given that $\widetilde{u}\leq\widetilde{u}_{-}$, we conclude $\widetilde{u}(x_{0})=\widetilde{u}_{-}(x_{0})$. Recall from Remark 2.9 that
$\widetilde{u}\leq T^{-}[\widetilde{u}]-\bar{S}\leq\cdots\leq\widetilde{u}_{-}.$
Thus, $T^{-}[\widetilde{u}](x_{0})=\widetilde{u}(x_{0})+\bar{S}$. Applying (c) of Remark 2.4, we find $x_{-1}\in\mathbb{R}^{d}$ such that
$T^{-}[\widetilde{u}](x_{0})=\widetilde{u}(x_{-1})+S(x_{-1},x_{0})$
This leads to
$\widetilde{u}(x_{0})+\bar{S}=T^{-}[\widetilde{u}](x_{0})=\widetilde{u}(x_{-1})+S(x_{-1},x_{0}).$
Therefore, $(x_{-1},x_{0})\in\mathcal{NS}(\widetilde{u})$, implying $x_{0}\in\pi_{2}(\mathcal{NS}(\widetilde{u}))=\mathcal{A}$. With $p_{0}=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}\widetilde{u}_{-}(x_{0})$, we have $(x_{0},p_{0})\in\mathcal{A}^{*}$. Consequently, *$\mathcal{I}^{*}(\widetilde{u}_{-})\subset\mathcal{A}^{*}.$ We have proved that $\mathcal{A}^{*}\subset\mathcal{I}^{*}(u)$ for each discrete weak K.A.M. solution $u$. Therefore, $\mathcal{A}^{*}=\mathcal{I}^{*}(\widetilde{u}_{-}).$
(3). It follows from items (iii) and (vi) of Proposition 2.14 that $\mathcal{A}^{*}\subset\mathcal{I}^{*}(u)$ for each discrete weak K.A.M. solution $u$. Thus, the desired result can be immediately deduced from item (2). ∎
2.2. Twist maps of the $2$-dimensional annulus
We begin by introducing the notion of twist maps.
Definition 2.16.
  1. (1)
    $f$ is isotopic to identity ;
  2. (2)
  3. (3)
    $f$ twists verticals to the right: if $F:\mathbb{R}\times\mathbb{R}^{*}\righttoleftarrow$, denoted by $F(x,p)=(y,p^{\prime})$, is a lift of $f$ then for all $x\in\mathbb{R}$, then there exists $\varepsilon>0$ such that
Definition 2.17.
  1. (1)
    $S$ is diagonal-periodic:
    $\forall(x,y)\in\mathbb{R}^{2},\quad S(x+1,y+1)=S(x,y);$
  2. (2)
    there exists $\varepsilon>0$ such that
    $\frac{\partial^{2}S}{\partial x\partial y}S(x,y)<-\varepsilon,\quad\forall(x,y)\in\mathbb{R}\times\mathbb{R}.$
  3. (3)
Remark 2.18.
For any lift $F$ of a twist map $f$, there exists a unique generating function $S$, up to addition by a constant, that satisfies Definition 2.17 (see [MMS89, Appendix 1]).
One can then state Aubry & Le Daeron Non-crossing Lemma:
Lemma 2.19 ([ALD83]).
Let $f$ be a twist map and $S$ be a generating function. If $x_{1},x_{2},y_{1},y_{2}$ are four reals such that $(x_{1}-x_{2})(y_{1}-y_{2})<0$, then
$S(x_{1},y_{1})+S(x_{2},y_{2})>S(x_{1},y_{2})+S(x_{2},y_{1}).$
This Non-crossing Lemma can be obtained by integrating over a quadrangle and we omit its detailed proof.
@@ 2.2. Twist maps of the $2$-dimensional annulus
Let $f$ be a twist map. For any $c\in\mathbb{R}$, the function
$S_{c}(x,y):=S(x,y)+c(x-y)$
satisfies Hypothesis 1.
@@
Items (4) and (5) of Hypothesis 1 straightforwardly follow from items (1) and (2) of Definition 2.17 respectively. Since $S$ is of class $C^{2}$, we deduce that $S_{c}$ is of class $C^{1}$ and locally semiconcave (see [CS04, Proposition 2.1.2]). Remaining work is to demonstrate that $S_{c}$ is coercive. This conclusion follows from the superlinearity of the generating function $S$, expressed as
$\lim_{|x-y|\rightarrow+\infty}\frac{S(x,y)}{|x-y|}=+\infty.$
It follows immediately from a lemma due to MacKay, Meiss and Stark’s article [MMS89, Lemma A1.6]
Remark 2.21.
  1. (a)
    The function $S_{c}$ is referred to as the generating function at the cohomology class $c$ of $\mathbb{T}$. As proved in Proposition 2.20, all theories in Section 2.1 are applicable to $S_{c}$. For instance, consider any $c$ in the first cohomology group $H^{1}(\mathbb{T},\mathbb{R})$, which is isomorphic to $\mathbb{R}$, the Discrete Weak K.A.M. Theorem 2.7 defines the effective interaction $\bar{S_{c}}$. The function $\alpha(c)=-\bar{S_{c}}$ is called Mather’s $\alpha$-function, which is $C^{1}$ (see [Zav23, Theorem 5.4.9]). As a second example, the Aubry set $\mathcal{A}_{c}^{*}$ at cohomology class $c$ is $F_{c}$-invariant, where
    $F_{c}(x,p)=F(x,p+c)-(0,c).$
  2. (b)
[…]
3. Proof of Theorem 1.1
3.1. $\Sigma_{+}$ is well-defined, non-decreasing and continuous
Let $u\in C^{0}(\mathbb{T})$. Define the set-valued map
$\Sigma_{+}:\mathbb{R}$ $\rightarrow 2^{\mathbb{R}}$
$x$$\mapsto\{y\in\mathbb{R}\mid T^{+}[u](x)=u(y)-S(x,y)\}.$
Thanks to Remark 2.4, we have $\Sigma_{+}(x)$ is non-empty for each $x\in\mathbb{R}$. For two subsets $A\subset\mathbb{R}$ and $B\subset\mathbb{R}$, we define $A\leq B$ if and only if $x_{1}\leq x_{2},\forall x_{1}\in A,x_{2}\in B.$

### v2

@@ 2.1. A notion of important invariant set
Suppose that $x\in\Sigma_{-}[v](y_{1})\cap\Sigma_{-}[v](y_{2})$. By item (a) of Lemma 2.19, we have
$\{-\partial_{1}S(x,y_{1}),-\partial_{1}S(x,y_{1})\}\subset\nabla^{-}v(x).$
Since $v$ is semiconcave, $\nabla^{+}v(x)$ is nonempty. Then, by item (iii) of Proposition A.2, $v$ is differentiable at $x$, and
$\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x)=-\partial_{1}S(x,y_{1})=-\partial_{1}S(x,y_{2}).$
Since the map $y\mapsto\partial_{1}S(x,y)$ is a homeomorphism (see item (2) of Hypothesis 2), it follows that $y_{1}=y_{2}$. ∎
Lemma 2.21.
If $v\in SC(\mathbb{T}^{d})$, then $v$ is $C^{1,1}$ on $\mathcal{O}[v]$. In other words, $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x)$ exists for all $x\in\mathcal{O}[v]$ and there exists $K>0$ such that
$\|\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x_{2})\|\leq K\|x_{1}-x_{2}\|,\quad\forall x_{1},x_{2}\in\mathcal{O}[v].$
Moreover, $T^{+}\circ T^{-}[v]$ takes the same values and derivatives as $v$ (so is also $C^{1,1}$) on $\mathcal{O}[v]$.
Proof.
For any $x\in\mathcal{O}[v]$, we have
(3) $T^{+}\circ T^{-}[v](x)=v(x).$
Since $T^{+}\circ T^{-}[v]\leq v$, item(ii) of Proposition A.2 implies that
$\nabla^{+}T^{+}\circ T^{-}[v](x)\supset\nabla^{+}v(x)\quad\text{and}\quad\nabla^{-}T^{+}\circ T^{-}[v](x)\subset\nabla^{-}v(x).$
Because $T^{+}\circ T^{-}[v]$ is semiconvex and $v$ is semiconcave, all four superdifferentials and subdifferentials are nonempty. Then by item (iii) of Proposition A.2, both $T^{+}\circ T^{-}[v]$ and $v$ are differentiable at $x$ and
(4) $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}\circ T^{-}[v](x)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x).$
Since $v$ is a semiconcave function, item (i) of Proposition A.2 guarantees the existence of a constant $C\geq 0$ such that
@@
$v(x_{1})-v(x_{2})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x_{2})(x_{1}-x_{2})\leq\frac{C}{2}|x_{1}-x_{2}|^{2},\quad\forall x_{1},x_{2}\in\mathcal{O}[v].$
On the other hand, because $T^{+}\circ T^{-}[v]$ is semiconvex, there exists $C^{\prime}\geq 0$ such that
$T^{+}\circ T^{-}[v](x_{1})-T^{+}\circ T^{-}[v](x_{2})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}\circ T^{-}[v]$ $(x_{2})(x_{1}-x_{2})$
$\geq-\frac{C^{\prime}}{2}\|x_{1}-x_{2}\|^{2},\quad\forall x_{1},x_{2}\in\mathcal{O}[v].$
Substituting equations (3) and (4), we replace $T^{+}\circ T^{-}[v]$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}\circ T^{-}[v]$ with $v$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v$ respectively. This gives
@@
$v(x_{1})-v(x_{2})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x_{2})(x_{1}-x_{2})\geq-\frac{C^{\prime}}{2}\|x_{1}-x_{2}\|^{2},\quad\forall x_{1},x_{2}\in\mathcal{O}[v].$
Combining these inequalities, we obtain
$-\frac{C^{\prime}}{2}\|x_{1}-x_{2}\|^{2}\leq v(x_{1})-v(x_{2})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x_{2})(x_{1}-x_{2})\leq\frac{C}{2}\|x_{1}-x_{2}\|^{2},\quad\forall x_{1},x_{2}\in\mathcal{O}[v].$
Exchanging the roles of $x_{1}$ and $x_{2}$, we get
$-\frac{C^{\prime}}{2}\|x_{1}-x_{2}\|^{2}\leq v(x_{2})-v(x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x_{1})(x_{2}-x_{1})\leq\frac{C}{2}\|x_{1}-x_{2}\|^{2},\quad\forall x_{1},x_{2}\in\mathcal{O}[v].$
Adding the two inequalities yields
$-C^{\prime}\|x_{1}-x_{2}\|^{2}\leq(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x_{2}))(x_{1}-x_{2})\leq C\|x_{1}-x_{2}\|^{2},\quad\forall x_{1},x_{2}\in\mathcal{O}[v].$
If we let $K=\max\{C,C^{\prime}\}$, then
$\|\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x_{2})\|\leq K\|x_{1}-x_{2}\|,\quad\forall x_{1},x_{2}\in\mathcal{O}[v].\qed$
Lemma 2.22.
Assume that $S$ satisfies Hypothesis 1 and Hypothesis 2. Let $u\in C^{0}(\mathbb{T}^{d})$.
  1. (a)
  2. (b)
Note that a version of Lemma 2.22 for discrete weak K.A.M. solutions could be referred in [Arn16].
Lemma 2.23.
If $u=T^{-}[v]$, $v\in SC(\mathbb{T}^{d})$ , then $\Sigma_{+}[u]$ is a continuous mapping on $\mathcal{O}[v]$. Moreover, if the discrete standard map $F$ is locally Lipschitz continuous, then $\Sigma_{+}[u]$ is Lipschitz continuous on $\mathcal{O}[v]$.
Proof.
For any $x\in\mathcal{O}[v]$, since $v$ is semiconcave, it follows from Lemma 2.21 that $T^{+}\circ T^{-}[v]=T^{+}[u]$ is differentiable at $x$. Then, by item (b) of Lemma 2.22, the set $\Sigma_{+}[u](x)$ is a singleton. Consequently, $\Sigma_{+}[u]$ can be viewed as a real-valued mapping on $\mathcal{O}[v]$.
For any $x\in\mathcal{O}[v]$, if $y=\Sigma_{+}[u](x)$, then, by (b) of Lemma 2.19 and the differentiability of $T^{+}[u]$ at $x$, we have $-\partial_{1}S(x,y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)$. From Lemma 2.21, which states $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[v](x)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x)$, it follows that $-\partial_{1}S(x,y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x)$. Using the definition of the discrete standard map $F$ (see Remark 1.9), we deduce that
$y=\pi_{1}\circ F(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x)).$
Thus, the continuity of $\Sigma_{+}[u]$ follows from the continuity of $F$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v$.
Now assume that $F$ is locally Lipschitz continuous, we will show that $\Sigma_{+}[u]$ is Lipschitz continuous on $\mathcal{O}[v]$. By item (v) of Proposition A.2, there exists $M>0$ such that $\|\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x)\|\leq M$ for any $x\in\mathcal{O}[v]$. Let $B_{M}:=\{p\in(\mathbb{R}^{d})^{*}\mid\|p\|\leq M\}$. Since $F$ is locally Lipschitz, the mapping $\pi_{1}\circ F$ is Lipschitz continuous on the compact set $\mathbb{T}^{d}\times B_{M}$. Denote the Lipschitz constant of $\pi_{1}\circ F$ by $K^{\prime}$. Then, for any $x_{1},x_{2}\in\mathcal{O}[v]$, we have
$\|\Sigma_{+}[u](x_{1})-\Sigma_{+}[u](x_{2})\|$ $=\left\|\pi_{1}\circ F(x_{1},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x_{1}))-\pi_{1}\circ F(x_{2},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x_{2}))\right\|$
$\leq K^{\prime}\|(x_{1},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{1}))-(x_{2},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2}))\|_{\mathbb{R}^{2d}}.$
Finally, by Lemma 2.21, we have
$\|(x_{1},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x_{1}))-(x_{2},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v(x_{2}))\|_{\mathbb{R}^{2d}}\leq\sqrt{1+K^{2}}\|x_{1}-x_{2}\|,$
where $K$ is the Lipschitz constant of $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}v$. Combining these, we conclude that
$\|\Sigma_{+}[u](x_{1})-\Sigma_{+}[u](x_{2})\|\leq K^{\prime}\sqrt{1+K^{2}}\|x_{1}-x_{2}\|.$
Thus, $\Sigma_{+}[u]$ is Lipschitz continuous on $\mathcal{O}[v]$. ∎
@@
Let $u$ be a discrete weak K.A.M. solution. Define $\mathcal{O}^{*}[u]:=\{(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x))\mid x\in\mathcal{O}^{\infty}[u]\}$. Then, it is an $F$-invariant set, i.e.,
$F(\mathcal{O}^{*}[u])=\mathcal{O}^{*}[u].$
@@
Since $u$ is a discrete weak K.A.M. solution, $T^{-}[u]=u+\bar{S}$.
Step 1. Show $F(\mathcal{O}^{*}[u])\subset\mathcal{O}^{*}[u]$: For any $x\in\mathcal{O}^{\infty}[u]$, by Proposition 2.8, $x\in\Sigma_{-}[u](\mathcal{O}^{\infty}[u]).$ That is, there exists $y\in\mathcal{O}^{\infty}[u]$ such that $x\in\Sigma_{-}[u](y)$. By Lemma 2.5, $y\in\Sigma_{+}[u](x)$. Using Lemma 2.19, we have
$-\partial_{1}S(x,y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)\quad\text{and}\quad\partial_{2}S(x,y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y).$
Hence, $F(\mathcal{O}^{*}[u])\subset\mathcal{O}^{*}[u].$
Step 2. Show $\mathcal{O}^{*}[u]\subset F(\mathcal{O}^{*}[u])$: For any $x\in\mathcal{O}^{\infty}[u]$, let $w\in\Sigma_{-}[u](x)$. By Proposition 2.8, $w\in\mathcal{O}^{\infty}[u]$. By Lemma 2.5, $x\in\Sigma_{+}[u](w)$. Using Lemma 2.19, we have
$-\partial_{1}S(w,x)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(w)\quad\text{and}\quad\partial_{2}S(w,x)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x).$
Hence, $F(\mathcal{O}^{*}[u])\supset\mathcal{O}^{*}[u].$
Combining both inclusions, we conclude that $F(\mathcal{O}^{*}[u])=\mathcal{O}^{*}[u].$
3. Proofs of Theorems 1.1-1.2
In Section 3 , we operate under the assumption that the function $S$ satisfies Hypothesis 1, Hypothesis 2 and Hypothesis 3. The proof of Theorem 1.1 is divided into the following four subsections and then Theorem 1.2 follows immediately.
3.1. $\Sigma_{+}[u]$ is a non-decreasing continuous real-valued mapping
Let $u\in C^{0}(\mathbb{T})$. Recall that the optimal forward mapping is the closed multi-valued mapping
$\Sigma_{+}[u]:\mathbb{R}$ $\rightarrow\mathbb{R}$
$x$ $\mapsto\mathop{\arg\max}_{y\in\mathbb{R}}\{u(y)-S(x,y)\}.$
So, without verification, $\Sigma_{+}[u](x)$ is a subset of $\mathbb{R}$. However, we will later show that under certain assumptions, $\Sigma_{+}[u](x)$ is actually a singleton for any $x\in\mathbb{R}$ (see Lemma 3.2). In other words, the optimal forward mapping is actually a mapping under certain assumptions.
Thanks to Remark 2.4 (a), we have $\Sigma_{+}[u](x)$ is non-empty for each $x\in\mathbb{R}$. For two subsets $A\subset\mathbb{R}$ and $B\subset\mathbb{R}$, we define $A\leq B$ if and only if $x_{1}\leq x_{2},\forall x_{1}\in A,x_{2}\in B.$